# SDEs with supercritical distributional drifts

### v1

@@ 2.2. Schauder’s estimate for heat equation in Besov spaces
$P_{t}f=p_{t}*f\quad\text{where}\quad p_{t}(x)=(4\pi t)^{-\frac{d}{2}}\mathrm{e}^{-\frac{|x|^{2}}{4t}}.$
@@
For any $T>0$, there is a constant $C=C(T,\alpha,d,p,q,p^{\prime},q^{\prime})>0$ such that for all $\lambda\geqslant 0$,
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}}_{p^{\prime},q}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}},$ (2.20)
@@
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}-\theta}_{p^{\prime},1}}\lesssim_{C,\theta}(1+\lambda)^{-\frac{\theta}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\infty}}.$ (2.21)
@@
Let $r\in[1,\infty]$ be defined by $1+\frac{1}{p^{\prime}}=\frac{1}{r}+\frac{1}{p}$ and $T>0$. For any $j\geqslant-1$, by Young’s inequality and Bernstein’s inequality (see [1, Lemma 2.1, p52]), one sees that
$\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p^{\prime}}$ $\leqslant\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}(p_{t-s}*f_{s})\|_{p^{\prime}}{\mathord{{\rm d}}}s$
=(2.3)${{=}}\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}p_{t-s}*\widetilde{\mathcal{R}}_{j}f_{s}\|_{p^{\prime}}{\mathord{{\rm d}}}s$
$\leqslant\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}p_{t-s}\|_{r}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}{\mathord{{\rm d}}}s$
$\lesssim 2^{(d-\frac{d}{r})j}\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}p_{t-s}\|_{1}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}{\mathord{{\rm d}}}s.$
Let $r^{\prime}\in[1,\infty]$ be defined by $1+\frac{1}{q^{\prime}}=\frac{1}{r^{\prime}}+\frac{1}{q}$. By Young’s inequality for time variable, we have
$\left(\int_{0}^{T}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p^{\prime}}^{q^{\prime}}{\mathord{{\rm d}}}t\right)^{\frac{1}{q^{\prime}}}\lesssim 2^{(d-\frac{d}{r})j}\left(\int_{0}^{T}\mathrm{e}^{-r^{\prime}\lambda s}\|{\mathcal{R}}_{j}p_{s}\|_{1}^{r^{\prime}}{\mathord{{\rm d}}}s\right)^{\frac{1}{r^{\prime}}}\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}}.$
@@
$\|{\mathcal{R}}_{j}p_{s}\|_{1}\lesssim_{C}1\wedge(2^{2j}s)^{-\theta},\ \ j\geqslant-1,\ s\in(0,T].$
If $r^{\prime}<\infty$, then by the change of variable, we have
$\left(\int_{0}^{T}\mathrm{e}^{-r^{\prime}\lambda s}\|{\mathcal{R}}_{j}p_{s}\|_{1}^{r^{\prime}}{\mathord{{\rm d}}}s\right)^{\frac{1}{r^{\prime}}}$ $\lesssim\left(\int_{0}^{\infty}\mathrm{e}^{-r^{\prime}\lambda s}[1\wedge(2^{2j}s)^{-2}]{\mathord{{\rm d}}}s\right)^{\frac{1}{r^{\prime}}}$
$=\left(2^{-2j}\int_{0}^{\infty}\mathrm{e}^{-r^{\prime}2^{-2j}\lambda s}[1\wedge s^{-2}]{\mathord{{\rm d}}}s\right)^{\frac{1}{r^{\prime}}}$
$\leqslant\left(2^{-2j}(1\wedge(r^{\prime}2^{-2j}\lambda)^{-1})\right)^{\frac{1}{r^{\prime}}}.$
If $r^{\prime}=\infty$, then we directly have
$\sup_{j\geqslant-1}\sup_{s\geqslant 0}\mathrm{e}^{-\lambda s}\|{\mathcal{R}}_{j}p_{s}\|_{1}\lesssim 1.$
@@
$\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}L^{p^{\prime}}}=\left(\int_{0}^{T}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p^{\prime}}^{q^{\prime}}{\mathord{{\rm d}}}t\right)^{\frac{1}{q^{\prime}}}$
$\qquad\lesssim 2^{(d-\frac{d}{r}-\frac{2}{r^{\prime}})j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r^{\prime}}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}}$
$\qquad=2^{(\alpha^{\prime}-2)j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r^{\prime}}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}},$
where $\alpha^{\prime}=d-\frac{d}{r}-\frac{2}{r^{\prime}}=\tfrac{2}{q}-\tfrac{2}{q^{\prime}}+\tfrac{d}{p}-\tfrac{d}{p^{\prime}}$.
Now, for any $\gamma\in[1,q]$ and $\theta\geqslant 0$, by the definition of Besov norm and Minkowskii’s inequality,
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}-\theta}_{p^{\prime},\gamma}}$ $=\left\|\left(\sum_{j\geqslant-1}2^{\gamma(2+\alpha-\alpha^{\prime}-\theta)j}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|^{\gamma}_{p^{\prime}}\right)^{\frac{1}{\gamma}}\right\|_{{\mathbb{L}}^{q^{\prime}}_{T}}$
$\leqslant\left(\sum_{j\geqslant-1}2^{\gamma(2+\alpha-\alpha^{\prime}-\theta)j}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|^{\gamma}_{{\mathbb{L}}^{q^{\prime}}_{T}L^{p^{\prime}}}\right)^{\frac{1}{\gamma}}$
$\lesssim\left(\sum_{j\geqslant-1}2^{\gamma(\alpha-\theta)j}(1\wedge(2^{-2j}\lambda)^{-\frac{\gamma}{r^{\prime}}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{\gamma}{q}}\right)^{\frac{1}{\gamma}}$
$=\left(\sum_{j\geqslant-1}2^{-\gamma\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{\gamma}{r^{\prime}}})\left(\int_{0}^{T}2^{q\alpha j}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{\gamma}{q}}\right)^{\frac{1}{\gamma}}.$
In particular, for $\theta=0$, if we take $\gamma=q$, then we get (2.20). For $\theta>0$, if we take $\gamma=1$, then
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}-\theta}_{p^{\prime},1}}\lesssim\sum_{j\geqslant-1}2^{-\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r^{\prime}}})\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\infty}}.$
Note that for $\theta\in(0,\frac{2}{r^{\prime}})=(0,2+\frac{2}{q^{\prime}}-\frac{2}{q})$,
$\sum_{j\geqslant-1}2^{-\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r^{\prime}}})\lesssim_{C}1\wedge\lambda^{-\frac{\theta}{2}}\lesssim_{C}(1+\lambda)^{-\frac{\theta}{2}}.$
@@
It is noted that (2.21) does not hold for $\theta=0$. Indeed, by (2.7) and Fourier’s transform,
@@
where $\asymp$ means that both sides are comparable up to a constant. Let $1\leqslant q\leqslant q^{\prime}\leqslant\infty$ and $1\leqslant p\leqslant p^{\prime}\leqslant\infty$ and $\theta\in(0,2-\frac{2}{q})$. By (2.20), (2.21) and ${\mathbf{B}}^{\alpha}_{p,q}\hookrightarrow{\mathbf{B}}^{\alpha}_{p,\infty}$, we immediately have
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-(\frac{2}{q}-\frac{2}{q^{\prime}}+\frac{d}{p}-\frac{d}{p^{\prime}})}_{p^{\prime},q}}+(1+\lambda)^{\frac{\theta}{2}}\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha-(\frac{2}{q}+\frac{d}{p}-\frac{d}{p^{\prime}})-\theta}_{p^{\prime},1}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (2.22)
@@ 3. A study of PDEs with distribution drifts
In this section, with $\lambda\geqslant 0$, our objective is to investigate the solvability of the nonhomogeneous partial differential equation:
@@
where $b$ and $f$ are distributions. We will establish the well-posedness of PDE (3.1) in two cases: the subcritical case and the supercritical case, utilizing different methods. In the subcritical case, we employ the Duhamel formula, while in the supercritical case, we utilize the maximal principle and the energy method. Throughout this section we fix a terminal time $T>0$.

### v3

@@ 2.2. Schauder’s estimate for heat equation in Besov spaces
$P_{t}f=\varphi_{t}*f\quad\text{where}\quad\varphi_{t}(x)=(4\pi t)^{-\frac{d}{2}}\mathrm{e}^{-\frac{|x|^{2}}{4t}}.$
@@
For any $T>0$ and $\gamma\in[q,q^{\prime}]$, there is a constant $C=C(T,\alpha,d,p,q,p^{\prime},q^{\prime},\gamma)>0$ such that
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}}_{p^{\prime},\gamma}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\gamma}},\ \ \lambda\geqslant 0,$ (2.20)
@@
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-\alpha^{\prime}-\theta}_{p^{\prime},1}}\lesssim_{C,\theta}(1+\lambda)^{-\frac{\theta}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\infty}},\quad\text{$\lambda\geqslant 0$}.$ (2.21)
@@
By virtue of the embedding (2.8) and the equivalence (see, for instance, [1, Proposition 2.78]),
$\|(1-\Delta)^{\alpha/2}f\|_{{\mathbf{B}}^{0}_{p,q}}\asymp\|f\|_{{\mathbf{B}}^{\alpha}_{p,q}},$
we may, without loss of generality, restrict our proof of (2.20) and (2.21) to the case $p^{\prime}=p$ and $\alpha=0$.
For any $j\geqslant-1$, by Young’s inequality and Bernstein’s inequality (see [1, Lemma 2.1, p52]), one sees that
$\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p}$ $\leqslant\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}(\varphi_{t-s}*f_{s})\|_{p}{\mathord{{\rm d}}}s$
=(2.3)${{=}}\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}\varphi_{t-s}*\widetilde{\mathcal{R}}_{j}f_{s}\|_{p}{\mathord{{\rm d}}}s$
$\leqslant\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}\varphi_{t-s}\|_{1}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}{\mathord{{\rm d}}}s$
$\lesssim\int_{0}^{t}\mathrm{e}^{-\lambda(t-s)}\|{\mathcal{R}}_{j}\varphi_{t-s}\|_{1}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}{\mathord{{\rm d}}}s.$
Let $r\in[1,\infty]$ be defined by $1+\frac{1}{q^{\prime}}=\frac{1}{r}+\frac{1}{q}$. By Young’s inequality for time variable, we have
$\left(\int_{0}^{T}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p}^{q^{\prime}}{\mathord{{\rm d}}}t\right)^{\frac{1}{q^{\prime}}}\lesssim 2^{(d-\frac{d}{r})j}\left(\int_{0}^{T}\mathrm{e}^{-r^{\prime}\lambda s}\|{\mathcal{R}}_{j}\varphi_{s}\|_{1}^{r}{\mathord{{\rm d}}}s\right)^{\frac{1}{r}}\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}}.$
@@
$\|{\mathcal{R}}_{j}\varphi_{s}\|_{1}\leqslant C[1\wedge(2^{2j}s)^{-\theta}],\ \ j\geqslant-1,\ s\in(0,T].$
If $r<\infty$, then by the change of variable, we have
$\left(\int_{0}^{T}\mathrm{e}^{-r\lambda s}\|{\mathcal{R}}_{j}\varphi_{s}\|_{1}^{r}{\mathord{{\rm d}}}s\right)^{\frac{1}{r}}$ $\lesssim\left(\int_{0}^{\infty}\mathrm{e}^{-r\lambda s}[1\wedge(2^{2j}s)^{-2}]{\mathord{{\rm d}}}s\right)^{\frac{1}{r}}$
$=\left(2^{-2j}\int_{0}^{\infty}\mathrm{e}^{-r2^{-2j}\lambda s}[1\wedge s^{-2}]{\mathord{{\rm d}}}s\right)^{\frac{1}{r}}$
$\leqslant\left(2^{-2j}(1\wedge(r2^{-2j}\lambda)^{-1})\right)^{\frac{1}{r}}.$
If $r=\infty$, then we directly have
$\sup_{j\geqslant-1}\sup_{s\geqslant 0}\mathrm{e}^{-\lambda s}\|{\mathcal{R}}_{j}\varphi_{s}\|_{1}\lesssim 1.$
@@
$\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}L^{p}}=\left(\int_{0}^{T}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{t}(f)\|_{p}^{q^{\prime}}{\mathord{{\rm d}}}t\right)^{\frac{1}{q^{\prime}}}$
$\qquad\lesssim 2^{-\frac{2}{r}j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}}$
$\qquad=2^{(\alpha^{\prime}-2)j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{1}{q}},$
where $\alpha^{\prime}=2-\frac{2}{r}=\tfrac{2}{q}-\tfrac{2}{q^{\prime}}$.
Now, for any $\gamma\in[1,q^{\prime}]$ and $\theta\geqslant 0$, by the definition of Besov norm and Minkowskii’s inequality,
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2-\alpha^{\prime}-\theta}_{p,\gamma}}$ $=\left\|\left(\sum_{j\geqslant-1}2^{\gamma(2-\alpha^{\prime}-\theta)j}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|^{\gamma}_{p}\right)^{\frac{1}{\gamma}}\right\|_{{\mathbb{L}}^{q^{\prime}}_{T}}$
$\leqslant\left(\sum_{j\geqslant-1}2^{\gamma(2-\alpha^{\prime}-\theta)j}\|{\mathcal{R}}_{j}{\mathscr{I}}^{\lambda}_{\cdot}(f)\|^{\gamma}_{{\mathbb{L}}^{q^{\prime}}_{T}L^{p}}\right)^{\frac{1}{\gamma}}$
$\lesssim\left(\sum_{j\geqslant-1}2^{\gamma(\alpha-\theta)j}(1\wedge(2^{-2j}\lambda)^{-\frac{\gamma}{r}})\left(\int_{0}^{T}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{\gamma}{q}}\right)^{\frac{1}{\gamma}}$
$=\left(\sum_{j\geqslant-1}2^{-\gamma\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{\gamma}{r}})\left(\int_{0}^{T}2^{q\alpha j}\|\widetilde{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{\gamma}{q}}\right)^{\frac{1}{\gamma}}.$
In particular, for $\theta=0$ and $\gamma\in[q,q^{\prime}]$, by $\left(\sum_{j}|a_{j}|^{\gamma}\right)^{1/{\gamma}}\leqslant\left(\sum_{j}|a_{j}|^{q}\right)^{1/q}$, we have
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2-\alpha^{\prime}-\theta}_{p,\gamma}}\lesssim\left(\sum_{j\geqslant-1}\left(\int_{0}^{T}2^{q\alpha j}\|{{\mathcal{R}}}_{j}f_{s}\|_{p}^{q}{\mathord{{\rm d}}}s\right)^{\frac{\gamma}{q}}\right)^{\frac{1}{\gamma}}\leqslant\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,\gamma}}.$
For $\theta>0$, if we take $\gamma=1$, then
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2-\alpha^{\prime}-\theta}_{p,1}}\lesssim\sum_{j\geqslant-1}2^{-\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r}})\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{0}_{p,\infty}}.$
Note that for $\theta\in(0,\frac{2}{r})=(0,2+\frac{2}{q^{\prime}}-\frac{2}{q})$,
$\sum_{j\geqslant-1}2^{-\theta j}(1\wedge(2^{-2j}\lambda)^{-\frac{1}{r}})\lesssim 1\wedge\lambda^{-\frac{\theta}{2}}\lesssim(1+\lambda)^{-\frac{\theta}{2}}.$
@@
Note that (2.21) does not hold for $\theta=0$. Indeed, ${\mathbf{B}}^{2}_{2,1}\subsetneqq{\mathbf{B}}^{2}_{2,2}$, ${\mathbf{B}}^{0}_{2,2}\subsetneqq{\mathbf{B}}^{0}_{2,\infty}$, and by (2.7) and Fourier’s transform,
@@
where $\asymp$ means that both sides are comparable up to a constant.
Remark 2.10.
Let $1\leqslant q\leqslant q^{\prime}\leqslant\infty$ and $1\leqslant p\leqslant p^{\prime}\leqslant\infty$ and $\theta\in(0,2-\frac{2}{q})$. By (2.20), (2.21) and ${\mathbf{B}}^{\alpha}_{p,q}\hookrightarrow{\mathbf{B}}^{\alpha}_{p,\infty}$, we immediately have
$\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{q^{\prime}}_{T}{\mathbf{B}}^{2+\alpha-(\frac{2}{q}-\frac{2}{q^{\prime}}+\frac{d}{p}-\frac{d}{p^{\prime}})}_{p^{\prime},q}}+(1+\lambda)^{\frac{\theta}{2}}\|{\mathscr{I}}^{\lambda}(f)\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha-(\frac{2}{q}+\frac{d}{p}-\frac{d}{p^{\prime}})-\theta}_{p^{\prime},1}}\lesssim\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (2.22)
@@ 3. A study of PDEs with distribution drifts
In this section, our goal is to investigate the solvability of the nonhomogeneous PDE:
@@
where $\lambda\geqslant 0$, $b$ and $f$ are distributions. We will establish the well-posedness of PDE (3.1) in two cases: the subcritical case and the supercritical case, utilizing different methods. In the subcritical case, we employ the Duhamel formula, while in the supercritical case, we utilize the maximal principle and the energy method. Throughout this section we fix a terminal time $T>0$.